# Training Latent Diffusion Models with Interacting Particle Algorithms

### v1

@@ A.4 The Full Training Algorithm
To avoid flat minima and ill-conditioning, we use an adaptive version of Langevin dynamics similar to Li et al., (2016); Kim et al., (2020). In particular, we update the particle cloud by running:
$z_{0,t+1}^{m,n}$ $\leftarrow z_{0,t}^{m,n}+Nh_{z}G_{t}^{m,n}M_{t}^{m,n}+\sqrt{\frac{2h_{z}}{M}}(G_{t}^{m,n})^{1/2}W_{t}^{m,n},$ (56)
$M_{t}^{m,n}$ $=aM_{t-1}^{m,n}+{N}(1-a)\nabla_{z_{0}^{m,n}}\hat{\mathcal{L}}(\theta_{t},\phi_{t},z^{1:M,1:N}_{0,t})$ (57a)
$V_{t}^{m,n}$ $=bV_{t-1}^{m,n}+(1-b)\text{diag}\left(\left[{N}\nabla_{z_{0}^{m,n}}\hat{\mathcal{L}}(\theta_{t},\phi_{t},z^{1:M,1:N}_{0,t})\right]^{\otimes 2}\right)$ (57b)
$G_{t}^{m,n}$ $=(V_{t}^{m,n}+\epsilon I)^{-1/2}$ (57c)
Here $\epsilon>0$ is a positive constant to avoid numerical instabilities, $10<a,b<$ 1 are hyperparameters like in Adam (Kingma and Ba,, 2014) (we take $a=0.9,b=0.999$ following the default), and we used the notation $\text{diag}(v^{\otimes 2})$ to denote the diagonal matrix with the $(i,j)^{th}$ entry being $v_{i}^{2}\delta_{ij}$. In practice, we set the preconditioner on the noise in (56) to $G_{t}^{m,n}=I$ during the first epoch for numerical stability.
Additionally, using the KL annealing scheme discussed in Section 3.2, we modify the loss in Equation 18 with:
^$\hat{\mathcal{L}}(\theta_{t},\phi_{t},z^{1:M,1:N}_{0,t}):=\frac{1}{N|\mathcal{B}|}\sum_{(i,m)\in[N]\times\mathcal{B}}\left[\gamma_{t}\hat{\mathcal{L}}_{\text{simple}}(\theta_{t},z_{0,t,m}^{i})-\log p_{\phi}(x_{m}|z_{0,t,m}^{i})\right],$ (58)
A.4.1 Derivation of the KL annealing
@@
Thus, we only need to adjust the amount of noise injected to the in (17) by setting $\sqrt{2h_{z}}$ to $\sqrt{2h_{z}\gamma_{t}}$ and weigh the diffusion loss $\mathcal{L}_{D}(\theta,z_{0})$ by $\gamma_{t}$.
@@
Algorithm 2 IPLD with Adaptive optimizer
1:Inputs: Training data points $\{x_{m}\}_{m\in[M]}$, optimizers $\texttt{Opt}_{\theta},\texttt{Opt}_{\phi}$, Step sizes $h_{\theta},h_{\phi},h_{z}$, KL weights $\gamma_{t}$, Initial particles $\{z_{m}^{i}\}_{(i,m)\in[N]\times[M]}$ sampled from $\mathcal{N}(0,I)$, Initial parameters $\phi,\theta$.
@@
5:     for $(i,m)\in[N]\times[M]$ do $\triangleright$ Update the particle cloud
@@ B.1 Details on the synthetic experiments
To train the models, we create the training data by first drawing 10K samples from: 1) a GMM with $25$ components of dimension $d_{z}=2$ as in Boys et al., (2024); Cardoso et al., (2024), and 2) a concentric circle distribution also of dimension $d_{z}=2$ similar to that in scikit-learn [^2] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_circles.html(Buitinck et al.,, 2013), then we project the data into a higher-dimensional ambient space with dimension $d_{x}=64$ using matrices $A\in\mathbb{R}^{d_{z}\times d_{x}}$ with orthogonal rows. We generate the matrices using the default implementation in PyTorch of orthogonal weight initialization (saxe2013exact). For better visualization, we set the first $2\times 2$ block to the identity for the concentric circle dataset.
@@
We use the AdamW optimizer (Loshchilov and Hutter,, 2019) with a learning rate of $1\times 10^{-3}$ for all models. For IPLD, we use a version of adaptive Langevin dynamics (Kim et al.,, 2020) resembling Adam (Kingma and Ba,, 2014) to optimize the particles; we set the step size to $1\times 10^{-1}$ for faster convergence. The KL annealing constant $c_{KL}$ is set to $0.01$ and number of warm up steps is set to $1000$ (cf. Section A.4).
@@ B.2 Details on the image experiments
We adopt the Diffusion Transformer architecture DiT-S, the smallest configuration, from Peebles and Xie, (2023) as the backbone for our latent diffusion model. We use a patch size of $1\times 1$, as our latent space of dimension $4\times 8\times 8$ is relatively small. For the Gaussian decoder $p_{\phi}(x|z_{0})$, we use a simplified version of VAE’s decoder without attention from (Rombach et al.,, 2022).
@@
Table 2: An overview of IPLD’s settings for the experiments. Here z-shape refers to the shape of the latent vector, which in the case of IPLD, is the shape of a single particle. We refer the readers to LDM (Rombach et al.,, 2022) and the implementation therein for more details on the architecture.
We comment that in line with recent discussions on latent diffusions (Dieleman,, 2025), the spatial structure of the latent space is crucial for the successful training of a latent diffusion model. Therefore, we choose to use a decoder that can induce explicit spatial structures in the latents $z$ instead of the ones used in the EBM literature (Yu et al.,, 2023), which compress the images to a single vector. However, we note our decoders have similar or smaller number of parameters than the implementation of Yu et al., (2023) (around 18M parameters).
@@
We use the exact same diffusion backbone and decoder for our re-implementation of VAE with diffusion prior (Wehenkel and Louppe,, 2021). For the encoder $q_{\phi}(z_{0}|x)$ of the VAE, we parametrize it with a Gaussian $\mathcal{N}(z_{0};\mu_{\phi}(x),\Sigma_{\phi})$, where $\Sigma_{\phi}$ is a diagonal matrix. The encoder’s architecture also follows from (Rombach et al.,, 2022), which an inverted version of the decoder.
@@ B.3 Warming-up with one particle
For the image experiments, we use a warm-started approach (Kuntz et al.,, 2023) for faster training. Namely, we initiate IPLD with a single particle $\{z_{m}\}_{m\in[M]}$ and run Algorithm 2 for 200 epochs on SVHN and CIFAR-10 (100 on CelebA64) before switching to $N>1$ particles by replicating the each particle $\{z_{m}\}_{m\in[M]}$ for $N$ times. We point out that due to the noise added during Langevin dynamics, the particles do not collapse to a single point.
@@ C.1 Interpolating the Latent Space
As a deep LVM, IPLD is able to learn a smooth latent space, enabling semantically meaningful interpolations. In Figure 4, we take two indices $m_{1},m_{2}\in[M]$ from the training set indices and extract the particles $z_{0}^{m_{1},0},z_{0}^{m_{2},0}$ from the particle cloud. We compute the linear interpolation via $\texttt{lerp}(x,y,s)=sx+(1-s)y$ between those particles for $s\in[0,1]$ and decode back into the image space with $g_{\phi}(z_{s,m_{1},m_{2}}^{(1)})$.
@@ C.2 Evolution of the Particle Cloud
We also visualize how the distribution $q(z_{0})$ evolves through the gradient flow. To achieve this, we save the particle cloud at training epochs $\{0,5,10,15,20,25,30,50,75,90\}$ and pass them through the decoder $g_{\phi}$ (Figures 5(a) and 5(b)). We note that there is an intriguing similarity between the particle evolution and the posterior mean $\mathbb{E}[x_{0}|x_{t}]$ of a diffusion model. For comparison, we show the evolution of the posterior mean predicted in Figures 6(a) and 6(b). The posterior mean is computed by a pre-trained score network taken from Song et al., 2021a in the pixel space via Tweedie’s formula (Efron,, 2011): $\mathbb{E}[x_{0}|x_{t}]\approx(\sqrt{\bar{\alpha}_{t}})^{-1}[x_{t}+\sigma_{t}^{2}\nabla_{x_{t}}{\bm{s}}_{\theta}(x_{t},t)]$, where $\sigma_{t}^{2}$ is the variance of the reverse process and ${\bm{s}}_{\theta}(x_{t},t)$ is the score estimating network. We remark that there have been several recent works attempting to delineate the connection between diffusion and gradient flow (Yi et al.,, 2023; Huang and Zhang,, 2023; Franceschi et al.,, 2023). Our method can be thought as an attempt of ’distilling’ the gradient flow into a diffusion model.
@@ C.3 Additional Uncurated Samples
Appendix D Proofs

### v2

@@ A.4 The Full Training Algorithm
We detail the algorithmic considerations in Section 3.3 and provide the pseudocode for practical training in Algorithm 2.
A.4.1 KL Divergence Annealing
@@
Thus, we only need to adjust the amount of noise injected to the in (17) by setting $\sqrt{2h_{z}}$ to $\sqrt{2h_{z}\gamma_{t}}$ and weigh the diffusion loss $\mathcal{L}_{D}(\theta,z_{0})$ by $\gamma_{t}$. Using the KL annealing scheme and re-weighted diffusion loss (40) discussed in Section 3.3, we thus modify the loss in (18) with:
^$\hat{\mathcal{L}}(\theta_{t},\phi_{t},z^{1:M,1:N}_{0,t}):=\frac{1}{N|\mathcal{B}|}\sum_{(m,n)\in\mathcal{B}\times[N]}\left[\gamma_{t}\hat{\mathcal{L}}_{\text{simple}}(\theta_{t},z_{0,t}^{m,n})-\log p_{\phi}(x_{m}|z_{0,t}^{m,n})\right],$ (62)
A.4.2 Preconditioning and Momentum
To avoid flat minima and ill-conditioning, we use an adaptive version of Langevin dynamics similar to Li et al., (2016); Kim et al., (2020) based on Adam (Kingma and Ba,, 2014). In particular, we update each particle by running:
$z_{0,t+1}^{m,n}$ $\leftarrow z_{0,t}^{m,n}+Nh_{z}G_{t}^{m,n}M_{t}^{m,n}+\sqrt{\frac{2h_{z}}{M}}(G_{t}^{m,n})^{1/2}W_{t}^{m,n},\quad\forall(m,n)\in[M]\times[N]$ (63)
$M_{t}^{m,n}$ $\leftarrow aM_{t-1}^{m,n}+{N}(1-a)\nabla_{z_{0}^{m,n}}\hat{\mathcal{L}}(\theta_{t},\phi_{t},z^{1:M,1:N}_{0,t})$ (64a)
$V_{t}^{m,n}$ $\leftarrow bV_{t-1}^{m,n}+(1-b)\text{diag}\left(\left[{N}\nabla_{z_{0}^{m,n}}\hat{\mathcal{L}}(\theta_{t},\phi_{t},z^{1:M,1:N}_{0,t})\right]^{\otimes 2}\right)$ (64b)
$G_{t}^{m,n}$ $\leftarrow(V_{t}^{m,n}+\epsilon I)^{-1/2}$ (64c)
Here $\hat{\mathcal{L}}$ is defined as in (62), $\epsilon>0$ is a positive constant to avoid numerical instabilities, $10<a,b<$ 1 are hyperparameters like in Adam (Kingma and Ba,, 2014) (we take $a=0.9,b=0.999$ following the default), and we used the notation $\text{diag}(v^{\otimes 2})$ to denote the diagonal matrix with the $(i,j)^{th}$ entry being $v_{i}^{2}\delta_{ij}$. In practice, we set the preconditioner on the noise in (63) to $G_{t}^{m,n}=I$ during the first epoch for numerical stability.
@@ A.4.2 Preconditioning and Momentum
Algorithm 2 IPLD with adaptive Langevin dynamics
1:Inputs: Training data points $\{x_{m}\}_{m\in[M]}$, optimizers $\texttt{Opt}_{\theta},\texttt{Opt}_{\phi}$, Step sizes $h_{\theta},h_{\phi},h_{z}$, KL weights $\gamma_{t}$, Initial particles $\{z_{0,0}^{m,n}\}_{(m,n)\in[M]\times[N]}$ sampled from $\mathcal{N}(0,I)$, Initial parameters $\phi,\theta$.
@@
5:     for $(m,n)\in\mathcal{B}\times[M]$ do $\triangleright$ Update the particle cloud
@@ B.1 Details on the synthetic experiments
We create the training data by first drawing 10,000 samples from: 1) a GMM with $25$ components of dimension $d_{z}=2$ as in Boys et al., (2024); Cardoso et al., (2024), and 2) a concentric circle distribution also of dimension $d_{z}=2$ similar to that in scikit-learn [^2] https://scikit-learn.org/stable/modules/generated/sklearn.datasets.make_circles.html(Buitinck et al.,, 2013), then we project the data into a higher-dimensional ambient space with dimension $d_{x}=64$ using matrices $A\in\mathbb{R}^{d_{z}\times d_{x}}$ with orthogonal rows. We generate the matrices using the default implementation in PyTorch of orthogonal weight initialization (Saxe et al.,, 2014). For better visualization, we set the first $2\times 2$ block to the identity for the concentric circle dataset.
@@
We train all models for 50 epochs with a batch size of 500. We use the AdamW optimizer (Loshchilov and Hutter,, 2019) with a learning rate of $1\times 10^{-3}$ for all models. For IPLD, we use a version of adaptive Langevin dynamics (Kim et al.,, 2020) (cf. Algorithm 2) to optimize the particles; we set the step size to $1\times 10^{-1}$ for faster convergence. The KL annealing constant $c_{KL}$ is set to $0.01$ and number of warm up steps is set to $1000$ (cf. Section A.4).
@@ B.2 Details on the image experiments
We adopt the Diffusion Transformer architecture DiT-S, the smallest configuration from Peebles and Xie, (2023), as the backbone for our latent diffusion model. We use a patch size of $1\times 1$, as our latent space of dimension $4\times 8\times 8$ is relatively small. For the Gaussian decoder $p_{\phi}(x|z_{0})$, we use a simplified version of VAE’s decoder without attention from Rombach et al., (2022).
@@
Table 2: An overview of IPLD’s settings for the experiments. Here z-shape refers to the shape of the latent vector, which in the case of IPLD, is the shape of a single particle. We refer the readers to LDM (Rombach et al.,, 2022) and the implementation thereof for more details on the architecture.
We comment that in line with recent discussions on latent diffusions (Dieleman,, 2025), the spatial structure of the latent space is crucial for the successful training of a latent diffusion model. Therefore, we choose to use a decoder that can induce explicit spatial structures in the latents $z$ instead of the ones used in the EBM literature (Yu et al.,, 2023), which compress the images to a flattened vector. However, we note our decoders have similar or smaller number of parameters than the implementation of Yu et al., (2023) (around 18 million parameters).
@@
We use the exact same diffusion backbone and decoder for our re-implementation of VAE with diffusion prior (Wehenkel and Louppe,, 2021). For the encoder $q_{\phi}(z_{0}|x)$ of the VAE, we parametrize it with a Gaussian $\mathcal{N}(z_{0};\mu_{\phi}(x),\Sigma_{\phi})$, where $\Sigma_{\phi}$ is a diagonal matrix. The encoder’s architecture also follows from Rombach et al., (2022), which an inverted version of the decoder.
@@ B.3 Warming-up with one particle
For the image experiments, we use a warm-started approach (Kuntz et al.,, 2023) for faster training. Namely, we initiate IPLD with a single particle $\{z_{m}\}_{m\in[M]}$ and run Algorithm 2 for 200 epochs on SVHN and CIFAR-10 (100 on CelebA64) before switching to $N>1$ particles by replicating the each particle $\{z_{m}\}_{m\in[M]}$ for $N$ times. We point out that due to the noise added in Langevin dynamics, the particles do not collapse to a single point.
@@ C.1 Interpolating the Latent Space
As a deep LVM, IPLD is able to learn a smooth latent space, enabling semantically meaningful interpolations. In Figure 4, we take two indices $m_{1},m_{2}\in[M]$ from the training set indices and extract the particles $z_{0}^{m_{1},0},z_{0}^{m_{2},0}$ from the particle cloud. We compute the linear interpolation via $\texttt{lerp}(x,y,s)=sx+(1-s)y$ between those particles for $s\in[0,1]$ and decode back into the image space with the decoder $g_{\phi}$.
@@ C.2 Evolution of the Particle Cloud
We also visualize how the particle cloud $z_{0}^{1:M,1:N}$ evolves through the gradient flow. To achieve this, we save the particle cloud at training epochs $\{0,5,10,15,20,25,30,50,75,90\}$ and pass them through the decoder $g_{\phi}$ (Figures 5(a) and 5(b)). We note that there is an intriguing perceptual similarity between the particle evolution and the posterior mean $\mathbb{E}[x_{0}|x_{t}]$ of a diffusion model. For comparison, we show the evolution of the posterior mean predicted in Figures 6(a) and 6(b). The posterior mean is computed by a pretrained score network taken from Song et al., 2021a in the pixel space via Tweedie’s formula (Efron,, 2011): $\mathbb{E}[x_{0}|x_{t}]\approx(\sqrt{\bar{\alpha}_{t}})^{-1}[x_{t}+\sigma_{t}^{2}\nabla_{x_{t}}{\bm{s}}_{\theta}(x_{t},t)]$, where $\sigma_{t}^{2}$ is the variance of the reverse process and ${\bm{s}}_{\theta}(x_{t},t)$ is the score estimating network. We remark that there have been several recent works attempting to delineate the connection between diffusion and gradient flow (Yi et al.,, 2023; Huang and Zhang,, 2023; Franceschi et al.,, 2023). Our method can be thought as an attempt of learning the gradient flow via a diffusion model.
@@ C.3 Additional Uncurated Samples
Appendix D Proofs of Theoretical Results